# Bi-DCSpell: A Bi-directional Detector-Corrector Interactive Framework for Chinese Spelling Check

### v1

@@ Abstract
Haiming Wu1, Hanqing Zhang1, Richeng Xuan2and Dawei Song1thanks: Corresponding authors 1School of Computer Science & Technology, Beijing Institute of Technology, Beijing, China 2Beijing Academy of Artificial Intelligence (BAAI), Beijing, China {wuhm, zhanghanqing, dwsong}@bit.edu.cn, rcxuan@baai.ac.cn
@@ 1 Introduction
Chinese Spelling Check (CSC) aims to automatically detect and correct spelling errors in Chinese sentences (Yu and Li, 2014; Wang et al., 2018; Huang et al., 2021). It serves as a foundation for a wide range of downstream applications in information retrieval (IR) and natural language processing (NLP), including Search Query Correction, Optical Character Recognition (OCR), and Essay Scoring. In recent years, CSC has garnered increasing attention from both academia and industry, emerging as a crucial task in the realm of Artificial Intelligence (Zhang et al., 2020; Li et al., 2024).
@@
Most existing approaches to CSC can be categorized into two types of framework: (a) correction-only (C-only) and (b) detection-to-correction (D2C). The C-only does not contain a detection module. Instead, it directly adopts a specially designed corrector to perform the correction task Hong et al. (2019); Zhang et al. (2020); Cheng et al. (2020); Li et al. (2021); Guo et al. (2021). The most recent C-only models are inspired by masked language model (MLM) Devlin et al. (2019), where each character to be corrected is predicted using contextual information. The second type of framework possesses both detection and correction modules but the interaction between them is single-directional in a way akin to a pipeline Cheng et al. (2020); Zhang et al. (2021); Zhu et al. (2022). That is, the detector first identifies the positions of errors, which are then used as prior knowledge for the correction module. While these two types of framework have been the dominant paradigms in the current CSC literature, they neglect the bi-directional interactions between detector and corrector. This oversight leads to issues such as the substitution of correct characters with incorrect ones (over-correction) or the failure to correct misspelled characters (under-correction).
Figure 1 presents an example to illustrate the differences among C-only, D2C, and bi-directional interactions. The input “{CJK}UTF8gbsn你的球机真厉害,我甘败下风。" contains two misspelled characters “{CJK}UTF8gbsn机(jī, machine)" and “{CJK}UTF8gbsn败(bài, fail)", which should be corrected as “{CJK}UTF8gbsn技(jì, skill)" and “{CJK}UTF8gbsn拜(bài, worship)", respectively. Figure 1(a) illustrates the C-only approaches that are typically underpinned by an MLM. Note that MLM tends to rectify a correct low-frequency collocation into a high-frequency one Yang and Yu (2022). In this example, as “{CJK}UTF8gbsn败(bài, fail)" is more frequent in natural language than “{CJK}UTF8gbsn拜(bài, worship)", it is difficult for a C-only approach to correct “{CJK}UTF8gbsn败(bài, fail)" properly in case there is no prior error information from detection. In Figure 1(b), the D2C interaction first uses an error detector that determines “{CJK}UTF8gbsn败(bài, fail)" is wrong. The result is then fed into the correction module, which then successfully corrects “{CJK}UTF8gbsn败(bài, fail)" as “{CJK}UTF8gbsn拜(bài, worship)". However, the corrector still may not be able to correct the other misspelled character “{CJK}UTF8gbsn机(jī, machine)" as it is not pre-detected by the detector. This reflects that one-directional interaction (from detection to correction) can still lead to under-correction or over-correction.
@@ 2.1 Correction-only Approaches for CSC
Correction-only (C-only) is a kind of approaches in CSC, employing a corrector to directly correct the input sequence without error detection. Recent approaches along this line with a masked language model (MLM), and have achieved significant performance improvements Hong et al. (2019); Wang et al. (2019); Zhang et al. (2020, 2021); Huang et al. (2021); Zhu et al. (2022). For instance, SpellGCN Cheng et al. (2020) integrates phonological and visual similarity information into character classifiers using a graph network, which then feeds the graph representation into MLM. ReaLise Xu et al. (2021) employs three distinct feature networks to capture phonetic, graphemic and semantic features, ultimately passing the fused representation through MLM. These methods focus on constructing CSC features and feeding them into an MLM-based corrector. In contrast, FASPell Hong et al. (2019) leverages phonological and visual similarity features to construct a filtering model, selects the most suitable candidate Chinese characters from a pre-trained Language Model (PLM). To better formalize this type of CSC methods, we denote the model parameters as $\bm{\theta}_{c}$, and the correction inference degenerates to:
@@ 2.2 Detection-to-correction for CSC
Detection-to-correction (D2C) represents the second type of existing CSC approaches, which first use an error detector to detect the position information of misspelled characters and then feed the detection results to the corrector. Most approaches employ a conventional multiple-stage pipeline approach, such as Soft-masked BERT Zhang et al. (2020), MLM-phonetics Zhang et al. (2021) and DR-CSC Huang et al. (2023). Unlike these pipeline D2C approaches, MDCSpell Zhu et al. (2022) uses parallel detection and correction feature representation modules, and the corrector receives the detector’s hidden states, thus, the inference in correction incorporates the feature from both detection and correction. Accordingly, D2C approaches can be formalized as:
@@ 3.2.2 Interactive Learning
To model the bi-directional interactions between detection and correction tasks, we design an interactive learning module as shown in the middle part of Figure 2. It consists of a series of stacked bi-directional interaction layers similar to the Transformer Vaswani et al. (2017), with each layer incorporating two task-specific cross-attention networks, two learnable control gates and a feed-forward network.
Task-specific Cross-attention Network. Firstly, two task-specific cross-attention networks use the feature knowledge learnt from one task to inform the representation learning for the other, respectively. Here, we use the same structure as vanilla attention Vaswani et al. (2017) in each cross-attention process, but the input involves two representation states (i.e., detection and correction) instead of one.
@@ 3.2.4 Model Training
where, the $\lambda\in[0,1]$ is a transfer-balanced hyperparameter to guide model learning.
@@ 4.1 Datasets and Pre-Processing
Extensive empirical evaluation is carried out on three widely used CSC datasets: SIGHAN13 Wu et al. (2013), SIGHAN14 Yu et al. (2014) and SIGHAN15 Tseng et al. (2015). We train Bi-DCSpell using four datasets, three of which are training data sets from SIGHAN, with approximately 10,000 data samples. The fourth dataset is an additional set of training data generated by an automatic method (Wang et al., 2018) with 271,009 samples.
Consistent with the current literature Xu et al. (2021); Zhu et al. (2022), we use three test datasets from SIGHAN13, SIGHAN14 and SIGHAN15 to evaluate the trained model. Furthermore, we use the same pre-processing procedure and transform the characters in these datasets to simplified Chinese with OpenCC [^1] https://github.com/BYVoid/OpenCC.
@@
SIGHAN13 Soft-Masked BERT Zhang et al. (2020) BERT 81.1 75.7 78.3 75.1 70.1 72.5
SpellGCN Cheng et al. (2020) BERT 80.1 74.4 77.2 78.3 72.7 75.4
MLM-phonetics Huang et al. (2021) BERT 82.0 78.3 80.1 79.5 77.0 78.2
DCN Wang et al. (2021) BERT 86.8 79.6 83.0 84.7 77.7 81.0
GAD Guo et al. (2021) BERT 85.7 79.5 82.5 84.9 78.7 81.6
MDCSpell Zhu et al. (2022) BERT 89.2 78.3 83.4 87.5 76.8 81.8
DORM Liang et al. (2023) ChineseBERT 87.9 83.7 85.8 86.8 82.7 84.7
DR-CSC Huang et al. (2023) ChineseBERT 88.5 83.7 86.0 87.7 83.0 85.3
@@
SIGHAN14 Soft-Masked BERT Zhang et al. (2020) BERT 65.2 70.4 67.7 63.7 68.7 66.1
SpellGCN Cheng et al. (2020) BERT 65.1 69.5 67.2 63.1 67.2 65.3
MLM-phonetics Zhang et al. (2021) BERT 66.2 73.8 69.8 64.2 73.8 68.7
DCN Wang et al. (2021) BERT 67.4 70.4 68.9 65.8 68.7 67.2
GAD Guo et al. (2021) BERT 66.6 71.8 69.1 65.0 70.1 67.5
MDCSpell Zhu et al. (2022) BERT 70.2 68.8 69.5 69.0 67.7 68.3
DORM Liang et al. (2023) ChineseBERT 69.5 73.1 71.2 68.4 71.9 70.1
DR-CSC Huang et al. (2023) ChineseBERT 70.2 73.3 71.7 69.3 72.3 70.7
@@
SIGHAN15 Soft-Masked BERT Zhang et al. (2020) BERT 67.6 78.7 72.7 63.4 73.9 68.3
SpellGCN Cheng et al. (2020) BERT 74.8 80.7 77.7 72.1 77.7 75.9
MLM-phonetics Huang et al. (2021) BERT 77.5 83.1 80.2 74.9 80.2 77.5
DCN Wang et al. (2021) BERT 77.1 80.9 79.0 74.5 78.2 76.3
GAD Guo et al. (2021) BERT 75.6 80.4 77.9 73.2 77.8 75.4
MDCSpell Zhu et al. (2022) BERT 80.8 80.6 80.7 78.4 78.2 78.3
CoSPA Yang and Yu (2022) ChineseBERT 79.0 82.4 80.7 76.7 80.0 78.3
DORM Liang et al. (2023) ChineseBERT 77.9 84.3 81.0 76.6 82.8 79.6
DR-CSC Huang et al. (2023) ChineseBERT 82.9 84.8 83.8 80.3 82.3 81.3
@@
Table 1: Experimental results on SIGHAN13, SIGHAN14 and SIGHAN15 test sets. Each model includes sentence-level precision, recall, and F1 score for both detection and correction. Due to the incompatibility of character-level results in the original paper Zhang et al. (2020), the results for Soft-Masked BERT in the table are sourced from Zhang et al. (2021), maintaining consistency in training data and metrics with our method. We evaluated Bi-DCSpell using two backbones, Chinese-BERT-wwm (BERT) and ChineseBERT for a fair comparison.
@@ 4.2 Model Settings
Considering that Bi-DCSpell is essentially a fine-tuning model for CSC, a wide range of CSC methods based on fine-tuning are selected as comparison models [^2] Note that, due to the inherent disparity between a fine-tuned model and one trained through a combination of pre-training and fine-tuning, we do not choose the models relying on pre-training for comparison, such as the current state-of-the-art CSC model PTCSpell Wei et al. (2023): PLM-FT is a fine-tuned PLM on the training data, which is indeed a standard C-only model and can be viewed as a simplified version of Bi-DCSpell without the interactive learning module and the detection encoder. Soft-Masked BERT Zhang et al. (2020), MLM-phonetics Zhang et al. (2021) and DR-CSC Huang et al. (2023) utilize a pipeline model, they all use the information of the previous step as the prior information of the next step. SpellGCN Cheng et al. (2020) incorporates phonological and visual similarity knowledge representation into BERT by employing a specialized graph convolutional network. DCN Wang et al. (2021) uses a dynamically connected network to measure the degree of dependence between any two adjacent Chinese characters. CoSPA Yang and Yu (2022) reports an alterable copy mechanism to increase the generation probability of the original input, thereby mitigating the over-correction. DORM Liang et al. (2023) allows the direct interaction between textual and phonetic information. MDCSpell Zhu et al. (2022) utilizes multi-task learning to jointly model detection and correction, implementing a multi-task learning model with task-specific feature encoders similar to those in Bi-DCSpell. However, it neglects the knowledge interaction between detection and correction.
@@ 4.4 Ablation Study
The range of optimal degree of interaction. The Bi-DCSpell framework contains two learnable controlling gates: $\alpha$ and $\beta$. We first use a heatmap to visualize the impact of these two gate values (set them as parameters) on the correction F1 score, as depicted in Figure 3. In general, a moderate bi-directional interaction yields the highest correction F1 score, outperforming both unidirectional ($\alpha=0.0$ or $\beta=0.0$) and fully bi-directional interactions ($\alpha=1.0$ and $\beta=1.0$). Specifically, the model achieves the optimal correction F1 score (darkest color) at ($\alpha=0.4$, $\beta=0.3$), with progressively lighter colors when moving away from this point. In other words, the F1 score gradually decreases, reaching a minimum at ($\alpha=0.0$, $\beta=0.0$) where there is no interaction.
@@ Appendix A Implementation Details
We use PyTorch Paszke et al. (2019) to implement the proposed Bi-DCSpell the Transformers library Wolf et al. (2020). For scientific comparison with existing methods, two pre-trained language models, Chinese-BERT-wwm Cui et al. (2021) (abbreviated as BERT) and ChineseBERT Sun et al. (2021), are used to initialize the embedding layer and the correction encoder, each model has 12 transformer layers with 12 attention heads and outputs for each token a hidden representation of dimensionality 768. We initialize the weights of the detector encoder using the Transformer parameters of the bottom two layers in the pre-trained language model. For model training, AdamW Loshchilov and Hutter (2018) is used as an optimizer with max epochs 20, the learning rate is set as 5e-5, and the batch size is set to 32. The training process takes about 10 hours on a single A6000 GPU (48GB GPU memory).

### v2

@@ Abstract
Anonymous ACL submission
@@ 1 Introduction
Chinese Spelling Check (CSC) aims to automatically detect and correct spelling errors in Chinese sentences (yu2014chinese; wang2018hybrid; huang2021phmospell). It serves as a foundation for a wide range of downstream applications in information retrieval (IR) and natural language processing (NLP), including Search Query Correction, Optical Character Recognition (OCR), and Essay Scoring. In recent years, CSC has garnered increasing attention from both academia and industry, emerging as a crucial task in the realm of Artificial Intelligence (zhang2020spelling; li2024mcrspell).
Most existing approaches to CSC can be categorized into two types of framework: (a) correction-only (C-only) and (b) detection-to-correction (D2C). The C-only does not contain a detection module. Instead, it directly adopts a specially designed corrector to perform the correction task hong2019faspell; zhang2020spelling; cheng2020spellgcn; li2021dcspell; guo-etal-2021-global. The most recent C-only models are inspired by masked language model (MLM) devlin2019bert, where each character to be corrected is predicted using contextual information. The second type of framework possesses both detection and correction modules but the interaction between them is single-directional in a way akin to a pipeline cheng2020spellgcn; zhang2021correcting; zhu2022mdcspell. That is, the detector first identifies the positions of errors, which are then used as prior knowledge for the correction module. While these two types of framework have been the dominant paradigms in the current CSC literature, they neglect the bi-directional interactions between detector and corrector. This oversight leads to issues such as the substitution of correct characters with incorrect ones (over-correction) or the failure to correct misspelled characters (under-correction).
@@
Figure 1 presents an example to illustrate the differences among C-only, D2C, and bi-directional interactions. The input “{CJK}UTF8gbsn你的球机真厉害,我甘败下风。" contains two misspelled characters “{CJK}UTF8gbsn机(jī, machine)" and “{CJK}UTF8gbsn败(bài, fail)", which should be corrected as “{CJK}UTF8gbsn技(jì, skill)" and “{CJK}UTF8gbsn拜(bài, worship)", respectively. Figure 1(a) illustrates the C-only approaches that are typically underpinned by an MLM. Note that MLM tends to rectify a correct low-frequency collocation into a high-frequency one yang2022cospa. In this example, as “{CJK}UTF8gbsn败(bài, fail)" is more frequent in natural language than “{CJK}UTF8gbsn拜(bài, worship)", it is difficult for a C-only approach to correct “{CJK}UTF8gbsn败(bài, fail)" properly in case there is no prior error information from detection. In Figure 1(b), the D2C interaction first uses an error detector that determines “{CJK}UTF8gbsn败(bài, fail)" is wrong. The result is then fed into the correction module, which then successfully corrects “{CJK}UTF8gbsn败(bài, fail)" as “{CJK}UTF8gbsn拜(bài, worship)". However, the corrector still may not be able to correct the other misspelled character “{CJK}UTF8gbsn机(jī, machine)" as it is not pre-detected by the detector. This reflects that one-directional interaction (from detection to correction) can still lead to under-correction or over-correction.
@@ 2.1 Correction-only Approaches for CSC
Correction-only (C-only) is a kind of approaches in CSC, employing a corrector to directly correct the input sequence without error detection. Recent approaches along this line with a masked language model (MLM), and have achieved significant performance improvements hong2019faspell; wang2019confusionset; zhang2020spelling; zhang2021correcting; huang2021phmospell; zhu2022mdcspell. For instance, SpellGCN cheng2020spellgcn integrates phonological and visual similarity information into character classifiers using a graph network, which then feeds the graph representation into MLM. ReaLise xu2021read employs three distinct feature networks to capture phonetic, graphemic and semantic features, ultimately passing the fused representation through MLM. These methods focus on constructing CSC features and feeding them into an MLM-based corrector. In contrast, FASPell hong2019faspell leverages phonological and visual similarity features to construct a filtering model, selects the most suitable candidate Chinese characters from a pre-trained Language Model (PLM). To better formalize this type of CSC methods, we denote the model parameters as $\bm{\theta}_{c}$, and the correction inference degenerates to:
@@ 2.2 Detection-to-correction for CSC
Detection-to-correction (D2C) represents the second type of existing CSC approaches, which first use an error detector to detect the position information of misspelled characters and then feed the detection results to the corrector. Most approaches employ a conventional multiple-stage pipeline approach, such as Soft-masked BERT zhang2020spelling, MLM-phonetics zhang2021correcting and DR-CSC huang2023frustratingly. Unlike these pipeline D2C approaches, MDCSpell zhu2022mdcspell uses parallel detection and correction feature representation modules, and the corrector receives the detector’s hidden states, thus, the inference in correction incorporates the feature from both detection and correction. Accordingly, D2C approaches can be formalized as:
@@ 3.2.2 Interactive Learning
To model the bi-directional interactions between detection and correction tasks, we design an interactive learning module as shown in the middle part of Figure 2. It consists of a series of stacked bi-directional interaction layers similar to the Transformer vaswani2017attention, with each layer incorporating two task-specific cross-attention networks, two learnable control gates and a feed-forward network.
Task-specific Cross-attention Network. Firstly, two task-specific cross-attention networks use the feature knowledge learnt from one task to inform the representation learning for the other, respectively. Here, we use the same structure as vanilla attention vaswani2017attention in each cross-attention process, but the input involves two representation states (i.e., detection and correction) instead of one.
@@ 3.2.4 Model Training
where the $\lambda\in[0,1]$ is a transfer-balanced hyperparameter to guide model learning.
@@ 4.1 Datasets and Pre-Processing
Extensive empirical evaluation is carried out on three widely used CSC datasets: SIGHAN13 wu2013chinese, SIGHAN14 yu2014overview and SIGHAN15 tseng2015introduction. We train Bi-DCSpell using four datasets, three of which are training data sets from SIGHAN, with approximately 10,000 data samples. The fourth dataset is an additional set of training data generated by an automatic method (wang2018hybrid) with 271,009 samples.
Consistent with the current literature xu2021read; zhu2022mdcspell, we use three test datasets from SIGHAN13, SIGHAN14 and SIGHAN15 to evaluate the trained model. Furthermore, we use the same pre-processing procedure and transform the characters in these datasets to simplified Chinese with OpenCC [^1] https://github.com/BYVoid/OpenCC.
@@
SIGHAN13 Soft-Masked BERT zhang2020spelling BERT 81.1 75.7 78.3 75.1 70.1 72.5
SpellGCN cheng2020spellgcn BERT 80.1 74.4 77.2 78.3 72.7 75.4
MLM-phonetics huang2021phmospell BERT 82.0 78.3 80.1 79.5 77.0 78.2
DCN wang-etal-2021-dynamic BERT 86.8 79.6 83.0 84.7 77.7 81.0
GAD guo-etal-2021-global BERT 85.7 79.5 82.5 84.9 78.7 81.6
MDCSpell zhu2022mdcspell BERT 89.2 78.3 83.4 87.5 76.8 81.8
DORM liang2023disentangled ChineseBERT 87.9 83.7 85.8 86.8 82.7 84.7
DR-CSC huang2023frustratingly ChineseBERT 88.5 83.7 86.0 87.7 83.0 85.3
@@
SIGHAN14 Soft-Masked BERT zhang2020spelling BERT 65.2 70.4 67.7 63.7 68.7 66.1
SpellGCN cheng2020spellgcn BERT 65.1 69.5 67.2 63.1 67.2 65.3
MLM-phonetics zhang2021correcting BERT 66.2 73.8 69.8 64.2 73.8 68.7
DCN wang-etal-2021-dynamic BERT 67.4 70.4 68.9 65.8 68.7 67.2
GAD guo-etal-2021-global BERT 66.6 71.8 69.1 65.0 70.1 67.5
MDCSpell zhu2022mdcspell BERT 70.2 68.8 69.5 69.0 67.7 68.3
DORM liang2023disentangled ChineseBERT 69.5 73.1 71.2 68.4 71.9 70.1
DR-CSC huang2023frustratingly ChineseBERT 70.2 73.3 71.7 69.3 72.3 70.7
@@
SIGHAN15 Soft-Masked BERT zhang2020spelling BERT 67.6 78.7 72.7 63.4 73.9 68.3
SpellGCN cheng2020spellgcn BERT 74.8 80.7 77.7 72.1 77.7 75.9
MLM-phonetics huang2021phmospell BERT 77.5 83.1 80.2 74.9 80.2 77.5
DCN wang-etal-2021-dynamic BERT 77.1 80.9 79.0 74.5 78.2 76.3
GAD guo-etal-2021-global BERT 75.6 80.4 77.9 73.2 77.8 75.4
MDCSpell zhu2022mdcspell BERT 80.8 80.6 80.7 78.4 78.2 78.3
CoSPA yang2022cospa ChineseBERT 79.0 82.4 80.7 76.7 80.0 78.3
DORM liang2023disentangled ChineseBERT 77.9 84.3 81.0 76.6 82.8 79.6
DR-CSC huang2023frustratingly ChineseBERT 82.9 84.8 83.8 80.3 82.3 81.3
@@
Table 1: Experimental results on SIGHAN13, SIGHAN14 and SIGHAN15 test sets. Each model includes sentence-level precision, recall, and F1 score for both detection and correction. Due to the incompatibility of character-level results in the original paper zhang2020spelling, the results for Soft-Masked BERT in the table are sourced from zhang2021correcting, maintaining consistency in training data and metrics with our method. We evaluated Bi-DCSpell using two backbones, Chinese-BERT-wwm (BERT) and ChineseBERT for a fair comparison.
@@ 4.2 Model Settings
Considering that Bi-DCSpell is essentially a fine-tuning model for CSC, a wide range of CSC methods based on fine-tuning are selected as comparison models [^2] Note that, due to the inherent disparity between a fine-tuned model and one trained through a combination of pre-training and fine-tuning, we do not choose the models relying on pre-training for comparison, such as the current state-of-the-art CSC model PTCSpell wei-etal-2023-ptcspell: PLM-FT is a fine-tuned PLM on the training data, which is indeed a standard C-only model and can be viewed as a simplified version of Bi-DCSpell without the interactive learning module and the detection encoder. Soft-Masked BERT zhang2020spelling, MLM-phonetics zhang2021correcting and DR-CSC huang2023frustratingly utilize a pipeline model, they all use the information of the previous step as the prior information of the next step. SpellGCN cheng2020spellgcn incorporates phonological and visual similarity knowledge representation into BERT by employing a specialized graph convolutional network. DCN wang-etal-2021-dynamic uses a dynamically connected network to measure the degree of dependence between any two adjacent Chinese characters. CoSPA yang2022cospa reports an alterable copy mechanism to increase the generation probability of the original input, thereby mitigating the over-correction. DORM liang2023disentangled allows the direct interaction between textual and phonetic information. MDCSpell zhu2022mdcspell utilizes multi-task learning to jointly model detection and correction, implementing a multi-task learning model with task-specific feature encoders similar to those in Bi-DCSpell. However, it neglects the knowledge interaction between detection and correction.
@@ 4.4 Ablation Study
The optimal degree of interaction. The Bi-DCSpell framework contains two learnable controlling gates: $\alpha$ and $\beta$. We first use a heatmap to visualize the impact of these two gate values (set them as parameters) on the correction F1 score, as depicted in Figure 3. In general, a moderate bi-directional interaction yields the highest correction F1 score, outperforming both unidirectional ($\alpha=0.0$ or $\beta=0.0$) and fully bi-directional interactions ($\alpha=1.0$ and $\beta=1.0$). Specifically, the model achieves the optimal correction F1 score (darkest color) at ($\alpha=0.4$, $\beta=0.3$), with progressively lighter colors when moving away from this point. In other words, the F1 score gradually decreases, reaching a minimum at ($\alpha=0.0$, $\beta=0.0$) where there is no interaction.
@@ Appendix A Implementation Details
We use PyTorch paszke2019pytorch to implement the proposed Bi-DCSpell the Transformers library wolf2020transformers. For scientific comparison with existing methods, two pre-trained language models, Chinese-BERT-wwm cui2021pre (abbreviated as BERT) and ChineseBERT sun2021chinesebert, are used to initialize the embedding layer and the correction encoder, each model has 12 transformer layers with 12 attention heads and outputs for each token a hidden representation of dimensionality 768. We initialize the weights of the detector encoder using the Transformer parameters of the bottom two layers in the pre-trained language model. For model training, AdamW loshchilov2018fixing is used as an optimizer with max epochs 20, the learning rate is set as 5e-5, and the batch size is set to 32. The training process takes about 10 hours on a single A6000 GPU (48Gb GPU memory).